OpenStudy (yourboydusty):

What is the distance between the two points graphed below? If necessary, round your answer to two decimal places.

OpenStudy (mathmale):

Use the distance formula, which you can find quickly through Google or in your textbook. Alternatively, note that the difference in the x-components is 6 and that the diff in the y-comps is 2. Using the Pythagorean Theorem, 6^2+2^2=d^2, where d is the distance between the 2 given points. Solve for d.

OpenStudy (yourboydusty):

?

OpenStudy (yourboydusty):

Can someone give me some help i dont understand this

OpenStudy (mathmale):

Glad to help, but I do expect that you do something up front to help yourself. Have you looked up "Pythagorean Theorem" and copied it down? Have you looked up "distance formula" and copied it down? If you'd please type out the distance formula here, I'll help you learn how to use it to calculate the distance between these two points. I have already given you an example of how you might use the Pythagorean Theorem here: "Alternatively, note that the difference in the x-components is 6 and that the diff in the y-comps is 2. Using the Pythagorean Theorem, 6^2+2^2=d^2, where d is the distance between the 2 given points. Solve for d." Please type out (or draw) the Distance Formula. You can easily find this formula by doing an Internet search.

OpenStudy (calculusxy):

Here is the distance formula: \(\large D = \sqrt{{(x_2 - x_1)} + {(y_2 - y_1})}\) This derives from the Pythagorean Theorem (\(\large a^2 + b^2 = c^2\)), specifically the equation that you would use to find the value of the hypotenuse. We know that \(a\) and \(b\) are lengths of the right triangle's legs.